OpenStudy (anonymous):

f(x) = 2(3)x, Section A is from x = 0 to x = 1 and Section B is from x = 2 to x = 3

OpenStudy (anonymous):

what is the average rate of change of each section

OpenStudy (anonymous):

ill fan and medal

OpenStudy (anonymous):

The average rate of change of each section is 5.

OpenStudy (anonymous):

so each section sepretly is .5

OpenStudy (anonymous):

@tb3_ty

OpenStudy (anonymous):

In Section A: Length of the section = (1 - 0) = 1 f(1) = 5 f(0) = 0 change in the value of the function = (5 - 0) = 5 Rate of change = (change in the value of the function) / (size of the section) = 5/1 = 5 In Section B: Length of the section = (3 - 2) = 1 f(3) = 15 f(2) = 10 change in the value of the function = (15 - 10) = 5 Rate of change = (change in the value of the function) / (size of the section) = 5/1 = 5

OpenStudy (anonymous):

/thank you

OpenStudy (anonymous):

welcome